Problem 5-22 CVP Applications; Contribution Margin Ratio; Break-Even Analysis; Cost Structure [LO5-1, LO5-3, LO5-4, LO5-5, LO5-6] Due to erratic sales of its sole product—a high-capacity battery for laptop computers—PEM, Inc., has been experiencing financial difficulty for some time. The company’s contribution format income statement for the most recent month is given below: Sales (13,000 units × $30 per unit) $ 390,000 Variable expenses 195,000 Contribution margin 195,000 Fixed expenses 217,500 Net operating loss $ (22,500 ) Required: 1. Compute the company’s CM ratio and its break-even point in unit sales and dollar sales. 2. The president believes that a $6,700 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $81,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the company’s monthly net operating income? 3. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $33,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)? 4. Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by $0.60 per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $4,800? 5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $51,000 each month. a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales. b. Assume that the company expects to sell 20,800 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.) c. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,800)?
Answer #1

1. Compute the company’s CM ratio and its break-even point in unit sales and dollar sales.

Variable cost per unit= $195000/13000= $15 per unit

Contribution margin per unit= Sales price per unit-Variable cost per unit

= $30-15= $15

Contribution margin ratio= Contribution margin*100/Sales

= $15*100/30= 50%

Break even point (unit) sales= Fixed cost/Contribution margin per unit

= $217500/15= 14500 units

Break even point dollar sales= Fixed cost/Contribution margin ratio

= $217500/50%= $435000

2. The president believes that a $6700 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $81,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the company’s monthly net operating income?

Increase in contribution margin ($81000*50%) $40500
Less: Increase fixed costs -6700
Increase in net operating income $33800

Net operating income will increases by $33800

3. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $33,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)?

New selling price= $30*0.90= $27 per unit

New sales units= 13000*2= 26000 units

New fixed costs= $217500+33000= $250500

Sales (26000 units*$27 per unit) $702000
Variable expenses (26000*$15) (390000)
Contribution margin 312000
Fixed expenses (250500)
Net operating income (loss) $61500

4. Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by $0.60 per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $4800?

New variable costs= $15+0.60= $15.60

New contribution margin per unit= Sales price per unit-Variable cost per unit

= $30-15.60= $14.40

Units= (Target profit+Fixed expenses)/Contribution margin per unit

= ($4800+217500)/14.40

= 15438 units

5. Refer to the original data. By automating, the company could reduce variable expenses by $3 per unit. However, fixed expenses would increase by $51000 each month.

a. Compute the new CM ratio and the new break-even point in unit sales and dollar sales.

New variable costs= $15-3=$12 per unit

New contribution margin per unit= Sales price per unit-Variable cost per unit

= $30-12= $18

New fixed expenses= $217500+51000= $268500

Contribution margin ratio= Contribution margin*100/Sales

= $18*100/30= 60%

Break even point (unit) sales= Fixed cost/Contribution margin per unit

= $268500/18= 14917 units

Break even point dollar sales= Fixed cost/Contribution margin ratio

= $268500/60%= $447500

b. Assume that the company expects to sell 20,800 units next month. Prepare two contribution format income statements, one assuming that operations are not automated and one assuming that they are. (Show data on a per unit and percentage basis, as well as in total, for each alternative.)

Not automated Automated
Total Per unit Percentage Total Per unit Percentage
Sales (20800*$30)= $624000 $30 100% (20800*$30)= $624000 $30 100%
Variable expenses (20800*$15)= -312000 15 (15*100/30)= 50% (20800*$12)= 249600 12 (12*100/30)= 40%
Contribution margin 312000 $15 50% 374400 $18 60%
Fixed expenses -217500 -268500
Net operating income $94500 $105900

c. Would you recommend that the company automate its operations (Assuming that the company expects to sell 20,800)?

Yes, the company should automate its operations as the net operating income of the company will increases if the company automated its operations